OpenStudy (anonymous):

Anyone know how to solve something like this? Show that: (27/64)^-2/3 = 9/16

OpenStudy (anonymous):

i'm pretty sure that 27/64 to the 2/3 power is the same is sqrt (27/64) to the third power

OpenStudy (anonymous):

\[\sqrt{27/64}^3\]

OpenStudy (anonymous):

to the negative third, rather.

myininaya (myininaya):

are you sure the other side isn't 16/9?

myininaya (myininaya):

(27/64)^(-2/3)=(64/27)^(2/3)=(4/3)^2=16/9